OpenStudy (anonymous):

Can some one graph 3 parabolas for me? :)

OpenStudy (anonymous):



OpenStudy (campbell_st):

go and download the webstart version of geogebra and it will do it for you. http://www.geogebra.org/cms/en/download/

OpenStudy (anonymous):

i CANNOT download anything on my laptop.

OpenStudy (anonymous):

I have the 3 equations, can someone just graph them real quick?

OpenStudy (anonymous):

Just choose a few points and find the vertices, intercepts, and a few more values. You'll have a fine graph.

OpenStudy (anonymous):

How do i find points from the equation?

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (campbell_st):

you choice any you line for a parabola try x = -3, -2, -1, 0, 1, 2, 3 or something like that

OpenStudy (anonymous):

Plug x values into the equation.

OpenStudy (anonymous):

Okay so if i plug in -2 as x i get the equation y=32-380+0 What do i do next?? @Numb3r1

OpenStudy (anonymous):

Solve for y and put that point on the graph.

OpenStudy (anonymous):

How do i solve for y? O.o

OpenStudy (anonymous):

she cannot graph taht

OpenStudy (anonymous):

If y=32-380, that's what y equals (just solve 32-380=?). However, @integralsabiti is probably right in saying that the point (-2, 380) will be well beyond the edge of your graph paper.

OpenStudy (anonymous):

I got -2,12

OpenStudy (anonymous):

What is the equation girl?

OpenStudy (anonymous):

Huh

OpenStudy (anonymous):

Give us an equation so we may help you. Or I'll give you an equation of parabola so you may graph it

OpenStudy (anonymous):

http://screencast.com/t/MVv0z5Kf